OpenStudy (anonymous):

how do i solve this?can you please explain to me step by step: ~ an initial charge of $20 to create the silk screen ~17.00/shirt for orders of 50 or fewer shirts ~15.80/shirt or orders of more than 50 shirts (all i have is $17 , x<= 50 and 15.50, x> 50, then i had no idea what to do with the 20) 1. write a piecewise function that gives the cost C for an order of x shirts 2.graph the function

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 20+17x & \text{if} & x \leq 50 \\ 20+15.8x& \text{if} & x > 50 \end{array} \right.\]

OpenStudy (anonymous):

how exactly did you get that?

OpenStudy (anonymous):

The fixed cost is $20 and the variable cost for the first 50 shirts is $17 time the number of shirts x. That is the first equation for cost.

OpenStudy (anonymous):

what mandolino said

OpenStudy (anonymous):

OHOHOH thanks!

OpenStudy (anonymous):

After 50 shirts the variable cost changes to $15.80 time the number of shirts.

OpenStudy (anonymous):

fixed cost it like the y - intercept of the line. even if you make no shirts you still have to pay $20

OpenStudy (anonymous):

ohhh,that's a big graph